Session: Feb 2003 (session 1)
Wavelength: 1.3 cm


To give feedback on this experiment, click here

Experiment Feedback

Medicina (Vargiu at Mon Feb 10 00:11:41 2003) Success :

Onsala (Michael Lindqvist at Mon Feb 10 10:06:10 2003) Success :
No known problems.

Noto (Pietro Cassaro at Mon Feb 10 11:40:40 2003) Minor failures :
In the final two hours of the experiment (from 22:00 UT) we had some problems with the correction of the subreflector position, so it is possible to have focus problems in the final scans.

Jodrell_Bank (Alastair Gunn at Mon Feb 10 14:47:20 2003) Success :
No known problems.

Cambridge (Alastair Gunn at Mon Feb 10 14:47:48 2003) Success :
No known problems. Please note that due to failure of headstack, track 30 has been remapped to track 34, i.e. third system track.

Metsahovi (Ari Mujunen at Wed Feb 12 12:10:53 2003) Success : RECO
Tracks 19,27,29,31,33 are probably quite bad, otherwise seemed to run okay.

Robledo (Z.P. (for Esther Moll) at Wed Feb 12 15:12:05 2003) Success :

Robledo (Z.P. (for Esther Moll) at Wed Feb 19 08:24:14 2003) Minor failures :
1) Level on video converters was stable during the whole experiment 2) Diode control worked for the whole activity 3) Full_cal procedure worked (of course only for LCP video converters : 1, 3, 5 and 7 ) 4) The formatter remained in sync. 5) During 4R pass the recorder#1 stopped at 6567 feet by itself. As a result scan 120 data from 15:19Z to 15:27Z was lost. Because of that pass 5F started from 6567 feet instead of 0 feet and the end of tape was reached at 15:42Z (scan 138 data from 15:42Z to 15:50Z was also lost). At 15:46Z the 'tapepos=17600' directive was entered and the tape was ready for scan 157. (See DR#M101405) 6) Sources of scans 157, 176 and 231 were lost due to the communication problem between EAC and the DSS63 antenna. > Antenna: DSS63 > Experiment name: ei005aro > Ops Crew: E/B > Weather condition: Clear > Station Time delay:-0.7897 > Recorder used: #1 > Number of Tapes: 2 > Label(/s):EVNT0820,E0A4 EVNT0934,10C5 > K-Band System Temp:53.3 > K-Band Y-factor:7.8